value: .0.25 points 3 attempts left Check my work A rocket can be powered by the reaction between dinitrogen tetroxide and hydrazine: Report pro Hint Solutio Guided Sol N204()+2 N2H4の→3N2(g) +4H2O(g) An engineer designed the rocket to hold 2.55 kg N,0, and excess gisg How much N2 would be produced according to the engineers design? Enter your answer in scientific notation. select)g N2 O Type here to search
Answer #1

N2O4(l) + 2N2H4(l) ---------------> 3N2(g) + 4H2O(g)

1 mole of N2O4 react with excess N2H4 to gives 3 moles of N2

92g of N2O4 react with excess N2H4 to gives 3*28g of N2

2550g of N2O4 react with excess N2H4 to gives = 3*28*2550/92   =2328.3g pf N2  

                                                                                                          = 2.3*10^3 g of N2 >>>>>>>answer